OpenStudy (anonymous):

find a third-degree polynomial equation with rational coefficients that has roots -4 and 6+i

OpenStudy (phi):

third-degree polynomial has 3 roots meaning x= 3 different values rational coefficients means the complex roots will come in complex conjugate pairs. They gave you one real root x= -4, and one complex root x= 6+i the third root must be the complex conjugate of 6+i the complex conjugate of a+bi is a- bi (negate the imaginary component)

OpenStudy (phi):

once you have all three roots: x= a, x=b and x=c you write each as (x-a)=0, (x-b)= 0 and (x-c) = 0 multiply them together to get your polynomial (x-a)(x-b)(x-c)